RawlsK29:

What is the value of z in the equation 5z − 9 = 36? 4 5 6 9

RawlsK29:

I dont understand how to work it out, everything I've tried just doesnt seem to work.

BlankSpace:

Let's solve your equation step-by-step. 5z−9=36 Step 1: Add 9 to both sides. 5z−9+9=36+9 5z=45

RawlsK29:

dont you have to do inverse operations

BlankSpace:

Step 2: Divide both sides by 5. 5z/5 = 45/5 z=9

RawlsK29:

I am still on step 1, wouldnt it be 5z + 9 - 9

BlankSpace:

Well, the way I do it, I add not subtract

RawlsK29:

Oh okay

BlankSpace:

So lets prove that the answer is correct

BlankSpace:

Plug it into the equation and in the place of the z

BlankSpace:

(5 x 9) - 9 = ?

BlankSpace:

What does 5 x 9 equal?

RawlsK29:

5(9) - 9 = 36 45 - 9 = 36 36 = 36

RawlsK29:

OMG, i would have never figured it out. Thank You!

BlankSpace:

No problem